### v1

# A Poisson-Guided Decomposition Network for Extreme Low-Light Image Enhancement

Abstract
Low-light image denoising and enhancement are challenging, especially when traditional noise assumptions, such as Gaussian noise, do not hold in majority. In many real-world scenarios, such as low-light imaging, noise is signal-dependent and is better represented as Poisson noise. In this work, we address the problem of denoising images degraded by Poisson noise under extreme low-light conditions. We introduce a light-weight deep learning-based method that integrates Retinex based decomposition with Poisson denoising into a unified encoder-decoder network. The model simultaneously enhances illumination and suppresses noise by incorporating a Poisson denoising loss to address signal-dependent noise. Without prior requirement for reflectance and illumination, the network learns an effective decomposition process while ensuring consistent reflectance and smooth illumination without causing any form of color distortion. The experimental results demonstrate the effectiveness and practicality of the proposed low-light illumination enhancement method. Our method significantly improves visibility and brightness in low-light conditions, while preserving image structure and color constancy under ambient illumination.
Index Terms:
Low-light imaging, image denoising, Poisson noise, retinex model, color distortion.
I Introduction
Images captured under low-light conditions often suffer from high noise levels, poor contrast, and distorted illumination, resulting in significant degradation in visual quality and make it difficult to extract meaningful information from images. This issue is prominently affecting real-time systems, such as autonomous driving [1], object detection [2] where accurate and clear image representation is essential for decision making. Therefore, efficient computational techniques are particularly important for reducing the loss of system performance caused by low-light images.
In recent decades, numerous conventional approaches for low-light image enhancement have been developed. Among them, traditional enhancement techniques such as histogram equalization [3, 4, 5] which enhances image contrast and brightness by redistributing the intensity levels of image pixels. These histogram equalization methods often have notable limitations, such as over-enhancement and noise amplification, particularly in extremely low-light conditions [6, 7]. Based on the Retinex model [8], researchers introduced single-scale Retinex [9] and multi-scale Retinex [10] for the low-light image enhancement. The core idea of the Retinex-based model is to decompose an image into reflectance and illumination components and then recover the illumination while preserving the reflectance, improving brightness and contrast [11, 12]. The authors in [13] proposed to enhance image illumination by estimating a structured illumination map [13]. Among the deep learning-based methods [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], RetinexNet [24], which enhances image contrast through brightness maps estimation and adjustment, with subsequent denoising using block matching and 3D filtering [25]. Ko et al. proposed a lightweight deep neural network model through knowledge distillation [26]. The authors in [27, 28] proposed vision transformer-based generative adversarial network to enhance low-light images.
Conventional image enhancement techniques such as [24, 29, 12, 30], [13, 31], often assume that noise is independent of the signal and is modeled as zero-mean additive Gaussian noise [32]. However, in practical scenarios such as low-light imaging the noise is typically signal-dependent and is better represented by Poisson statistics [33, 34]. The noise in extreme low-light images is dominated by Poisson statistics, which differs from the Gaussian noise assumption made by many standard denoising algorithms [35, 36]. This distinction makes conventional methods less effective in handling the noise typically found in low-light environments [37]. The illumination in such images is often distorted due to extreme low illumination, making it challenging to recover true scene details. As a result, the performance of most existing techniques is limited under low-light conditions, where the signal-to-noise ratio (SNR) is low and the photon counts are limited [38, 39, 40].
In this paper, we propose a deep learning method for illumination recovery and enhancement of images corrupted by Poisson noise due to poor lighting conditions. In particular, we introduce a unified encoder-decoder network which integrates Retinex based decomposition with Poisson denoising. Our encoder-decoder network learns an effective decomposition process while ensuring consistent reflectance and smooth illumination components. Our proposed modified Retinex decomposition has an additional component representing Poisson noise at each pixel. Our novelty is that we have a dedicated loss function in our deep learning architecture to suppress the noisy component in the output/enhanced image.
The rest of the paper is organized as follows. In Section II, we introduce the Poisson denoising problem and our proposed Poisson guided decomposition of extreme low-light image. We present our proposed deep learning network in Section III and experimental results in Section IV. Finally, we conclude with a summary in Section V.
Figure 1: Proposed architecture for Poisson noise removal and low-light image enhancement. The network is structured with a convolutional encoder-decoder framework. The three branches of the output layers generate an illumination map, a reflectance map, and a noise map.
II Poisson denoising problem
The primary source of noise in digital images arises during image acquisition or transmission. In the case of low-light images with poor exposure, noise is predominantly due to a low signal-to-noise ratio (SNR) and low photon count: fewer photons reaching the sensor limit the number of photons collected by a pixel. This results in increased uncertainty, as described by the Poisson distribution, which models the occurrence of independent random events such as photon arrivals. Consequently, Poisson noise, also known as shot noise, becomes the dominant form of noise in such conditions.
Let $X$ denote a noisy image produced by a sensor. The objective of denoising is to recover a clean latent image $Y$. In extreme low-light imaging, the noise is dominated by Poisson noise. Given the true value $Y(i)$ of the $i$-th pixel expressed in the number of photons, the corresponding value of the observed pixel $X(i)$ is an independent Poisson distributed random variable;
$P(X(i)=n)=\frac{\lambda^{n}e^{-\lambda}}{n!},$ (1)
where $\lambda>0$ is both mean and variance [41]. In defining the Poisson statistics for each pixel of the low-light image, it is assumed that the number of arrivals during non-overlapping time intervals are independent random variables and there exists a positive quantity $\lambda$ which is the contribution of a photon to the intensity of the pixel in $\Delta t$ duration and satisfying the postulates for the distribution of Poisson for each value of the pixels [41].
II-A Our Poisson-Guided Image Decomposition
It is well known that low-light digital images are typically captured and stored in RGB format [42], and since Poisson noise in the sensor affects all three color channels. This can lead to color distortion [43], [44] during image enhancement. To preserve color constancy under ambient illumination, we consider to decompose the captured color image $X$ as following:
$X^{c}(i)=L(i)\circ R^{c}(i)\circ N^{c}(i),\quad c\in\{r,g,b\},$ (2)
where $\circ$ indicates element-wise multiplication at each pixel $i$; $L$, $R$, and $N$ are illumination, reflectance, and noisy components respectively. Poisson noise is neither additive nor multiplicative, but it is signal-dependent. However, it can be formulated as a multiplicative component of the noise in the pixel values. To preserve color constancy and prevent color shifts during enhancement, we formulate a Poisson-guided image decomposition framework in which an image is decomposed into element-wise multiplication of three components: illumination $L$, reflectance ($R$) and noisy ($N$).
Figure 2: Visual comparison with other state-of-the-art low-light enhancement methods: (a) low-light input image, (b) LIME [13], (c) deep Retinex [24], (d) SRIE [45], (e) zero-DCE [46], (f) Proposed, (g) Ground Truth. Our method produces superior quality output image without any color distortion.
III Proposed Method
III-A Network Architecture
We propose a light-weight convolutional encoder-decoder network shown in Figure 1.
III-A1 Initial Convolution Layer
The first layer is a convolutional layer with 3 input channels (RGB input images patches) of size $3\times 128\times 128$ and $64$ output channels, utilizing a $3\times 3$ kernel size. This layer extracts basic low-level features from the input image, such as edges and textures.
III-A2 Encoder-Decoder Layers
The encoder consists of three consecutive convolutional layers with strides of two. Each convolutional layer is followed by a ReLU activation function to introduce nonlinearity for better feature representation. The decoder consists of three deconvolutional layers, each taking input from both the corresponding encoder layer (via concatenation) and the previous decoder layer. This enables the network to recover spatial resolution while combining high-level and low-level features for better image reconstruction.
III-A3 Image Reconstruction (Output Layers)
The network is divided into three parallel output layers to predict three distinct components: illumination map (via sigmoid activation function), reflectance map (via sigmoid activation) and noise (using a tanh activation function). Each output is resized to match the original input dimensions. Finally, the output enhanced image is the product of illumination map and reflectance map.
III-B Loss Function
The loss function is designed to optimize the performance of the model by incorporating multiple components that guide the network to perform image reconstruction, noise suppression, and smoothness regularization. Our proposed loss function is as follows:
$\mathcal{L}_{total}=\mathcal{L}_{rec}+\lambda_{1}\mathcal{L}_{ret}+\lambda_{2}\mathcal{L}_{n},$ (3)
where $\mathcal{L}_{rec}$, $\mathcal{L}_{ret}$, $\mathcal{L}_{n}$ denote reconstruction loss, Retinex estimation loss, and noise estimation loss respectively. The losses are weighted by hyper-parameters $\lambda_{1}$, $\lambda_{2}$.
III-B1 Reconstruction Loss
The error in decomposition is regulated by reconstruction loss function as follows:
$\mathcal{L}_{rec}=\parallel X-(R\circ L\circ N)\parallel_{1}$ (4)
which measures the fidelity between the original and reconstructed images while providing robustness to outliers and encouraging sparsity in the reconstruction error.
III-B2 Retinex Estimation Loss
This loss component is expressed as a combination of loss due to decomposition loss $\mathcal{L}_{decom}$ and smoothness preserving smoothness loss [47] $\mathcal{L}_{sps}$ follows,
$\mathcal{L}_{ret}=\gamma\mathcal{L}_{decom}+(1-\gamma)\mathcal{L}_{sps}$ (5)
where $\gamma$ is a hyperparameter. Further, loss due to illumination-reflectance decomposition error is given by:
$\mathcal{L}_{decom}=\parallel Y-R\circ L\parallel_{1}.$ (6)
The smoothness preserving smoothness loss is given by:
$\mathcal{L}_{sps}=\left\|\Delta L\circ\ exp{(-\beta\Delta R)}\right\|_{1}$ (7)
whereas, $\beta$ represents the coefficients for balanced structure and smoothness of illumination map, which is defined as $10$ in our experiments and $\Delta R$ represents gradient of reflectance in both x and y direction, $\Delta L$ represents gradient of Illumination in both x and y direction. The term exponential is used to protect against structural loss in the illumination map, and the gradient ensures the smoothness of the illumination.
III-B3 Noise Loss
The loss term works as a constraint to extract the noisy component from the low-light image:
$\mathcal{L}_{n}=\parallel N-\frac{\text{Poisson}(Y)}{Y+\alpha}\parallel_{1}$ (8)
which is used to suppress the intensity of noise ($\alpha$ is set as $10^{-6}$ computational stability).
Figure 3: Visual comparison with other state-of-the-art low-light enhancement methods: (a) low-light input image, (b) LIME [13], (c) deep Retinex [24], (d) SRIE [45], (e) zero-DCE [46], (f) Proposed, (g) Ground Truth. Notice that our method achieves better quality in the enhanced image than the other methods.
IV Experimental Results
IV-1 Dataset and Experimental Setup
We have used the Low Light dataset (LOL dataset) [24], which contains 500 paired low-light and normal-light images which are collected by changing exposure time and ISO, while other configurations of the cameras are fixed. During training (100 epochs), each input consists of $3\times 128\times 128$ sized RGB color patches of paired low-light and normal-light images.
IV-2 Baseline Methods
To demonstrate the effectiveness of our proposed methods, we compare with the following baseline methods: LIME [13], deep Retinex [24], SRIE [45], Light-weight low-light enhancement [26], and zero-DCE [46].
IV-3 Visual Comparison
We display the visual results of our method along with the baseline ones on two different scenes at varying illumination levels in Figures 2 and 3. In both cases, we have considered input image with four different levels of illumination (reduced as we move down from the first row to the fourth row). Notice that in both Figures 2 and 3, there are color distortions in the output images by the baseline methods from (b) to (f). Our method produces best visual quality restored images without color distortion shown in Fig. 2(f) and Fig. 3(f). Another important fact is that such color distortion gets worse with very low level of illumination (as is the case for last rows in both Figures 2 and 3) which is the case of extreme low-light condition. In such cases, our method significantly outperforms the baseline methods.
IV-4 Quantitative Comparison
We use three referenced and one no-referenced image metrics to quantify the comparative performance evaluation of our method with benchmarks: peak signal-to-noise ratio (PSNR) [48], structural similarity index measure (SSIM) [48], mean absolute error (MAE) [49], and no-reference image quality evaluator (NIQE) [48]. For PSNR and SSIM, a higher value indicates better quality, whereas for NIQE and MAE, a lower value indicates better quality of the enhanced image. In Table I, we summarize the performance of low-light enhancement using our method and the baselines. We found that our method outperforms the existing methods in terms of both PSNR and SSIM [50]. In addition, our method achieves competitive performance with the baseline methods in terms of NIQE and MAE metrics. Experimental results confirm that our method can perform efficient low-light enhancement even in extreme low-light conditions.
Our research demonstrates that the dominance of Poisson noise in extreme low-light images significantly contributed to color distortion, which gets severe with extreme low-lightness. Poisson noise, being signal-dependent, is more prominent in low-light conditions where photon counts ($c$) are low. Visual results confirm our method can perform efficient low-light enhancement even in extremely low-light conditions.
TABLE I: Comparison for different methods. The best are marked as red, the second best are marked as blue.
Method PSNR(dB) SSIM NIQE MAE
LIME [13] 15.66 0.42 9.35 35.57
Deep Retinex [24] 16.06 0.58 4.31 35.78
SRIE [45] 9.85 0.40 7.72 80.00
Zero-DCE [46] 12.35 0.55 3.42 61.71
Light-weight [26] 15.99 0.59 4.54 —-
Proposed 17.12 0.67 3.54 32.33
V Conclusion
In this work, we attempt to address a challenging task of extreme low-light image enhancement and denoising. We demonstrate with experimental finding that many baseline methods suffer from color distortion when the illumination level of the scene/object is extremely poor. We introduced a novel image decomposition framework in contrast to the Retinex model of image decomposition. Further, our encoder-decoder-based deep learning method could produce superior recovery of extreme low-light images without any visible color distortion. Given that Poisson noise dominates in extreme low-light images because of limited photon counts, specialized Poisson denoising techniques are essential for preserving image quality and structural details. In future research, we will focus on developing denoising models that can robustly handle mixed noise distributions while enabling more efficient solutions for real-time applications.